OpenStudy (anonymous):

z^2 + 15z + 56 = 0 What Are The Solutions? Please Help, I'd Appreciate it! =)

OpenStudy (anonymous):

factor \[z^2+15z+56\] \[z^2+15z+56=0\] \[(z+7)(z+8)=0\] \[z+8=0\] \[z=-8\] or \[z+7=0\] \[z=-7\]

OpenStudy (anonymous):

Thank youuuuuuuu(=

OpenStudy (anonymous):

@ wow you can't solve this..hahaha

OpenStudy (anonymous):

now now, no need to be a wingspan.

OpenStudy (anonymous):

hush, hush little piggy